OpenStudy (anonymous):

@zepdrix This one feels tricky - I don't know how to get the y's on one side...

OpenStudy (anonymous):

\[y' \cosh^2{x}=\sin^2{y}\] y(0) = pi/2

OpenStudy (amistre64):

division ....

zepdrix (zepdrix):

lol :)

OpenStudy (anonymous):

What I meant with getting them on one side, was getting the y's there and how to use them then haha I get \[\frac{1}{\sin^2y} \text{dy}=\frac{1}{\cosh^2y}\text{dx}\] but don't know how to integrate that

OpenStudy (anonymous):

the cosh^2 has an x, sorry, not a y

zepdrix (zepdrix):

Remember your trig identities? \(\large \dfrac{1}{\sin\theta}=?\)

zepdrix (zepdrix):

The right side looks a little tricky... I'd have to brush up on my hyperbolic identities... I'm assuming that gives us \(\large sech^2x\) though..

OpenStudy (amistre64):

arent the hyper trigs "e"-able?

zepdrix (zepdrix):

They're e-able.. but doesn't that put a bunch of e's in the denominator? I think that will be a little tricky.. hmm :O Maybe it's worth trying.

OpenStudy (anonymous):

Ahh! \[\csc^2y \text{ dy} = sech^2x\text{ dx}\]?

zepdrix (zepdrix):

Yah looks good so far :)

OpenStudy (amistre64):

these tables might help jog a memory cell or two ;) http://math2.org/math/integrals/tableof.htm

OpenStudy (amistre64):

... but then its squared :/

OpenStudy (amistre64):

i see that the h makes no difference. tan down to sec^2 tanh down to sech^2

zepdrix (zepdrix):

ya, good times.

OpenStudy (anonymous):

Sorry, I'm back. My Internet disconnected

OpenStudy (anonymous):

Okay, so integrating that part that I wrote, I get \[-\cot y = \tanh^2 x +c\] Now what?

zepdrix (zepdrix):

Oh we were given initial conditions? Hmm ok.\[\large \cot y=C-\tanh x\] So plug the value in and let's see what we get.\[\cot\left(\frac{\pi}{2}\right)=C-\tanh 0\]

OpenStudy (anonymous):

C = 0 so cot y = - tanh x and y = arccot(-tanh x) ?

zepdrix (zepdrix):

Yah that appears to be correct. What a weird problem lol.

OpenStudy (anonymous):

Yeah some of these problems are a little weird haha Thanks you guys for the help!